OpenStudy (krl1357):

If 78.4 grams of aluminum metal react with excess hydrochloric acid, how many grams of hydrogen gas can be produced?

OpenStudy (anonymous):

First, I think you need to write out a balanced equation. Al + HCl ----> H2....etc and then I think you write out the molar mass of each product and reactants

OpenStudy (michele_laino):

HInt: I think that such process can be described by the subsequent equation of reaction: \[{\text{Al}}\left( s \right) + 3{\text{HCl}}\left( l \right) \to {\text{AlC}}{{\text{l}}_3} + \frac{3}{2}{{\text{H}}_2}\left( g \right)\] so, for \(1\) mole of \(Al\), I get \(3/2\) mol of \(H_2\) gas